OpenStudy (anonymous):

Three brothers spend different amounts of money on their recent vacation. Matt spent $25 less than Todd. Jeff spent 3 times as much as matt. How much did each brother spend if they spent a total of $275

rishavraj (rishavraj):

let the amount tht Todd spent = x sp Matt spent = x + 25 and Jeff spent = 3(x + 25) and they spent a total of $275